Grade 10Mechanics

The relative density of a material is found by weighing the body first in air and then in water. If the weight in air
is (10.0 ± 0.1) gf and weight in water is (5.0 ± 0.1) gf, then the maximum permissible percentage error in
relative density is

aakanshika kapoor
11 Years agoGrade 10

1 Answer

Chaya
8 Years ago
Relative density=Weight in air/Loss of weight in waterLoss of weight in water=(10.0-5.0)+_(0.+0.1)=5.0+_0.2N% error in relative density=%error in weight of air+%error in loss of weight in water=(0.1/10.0+0.2/5.0)*100=5%